# DIRAS: Efficient LLM-Assisted Annotation of Document Relevance in Retrieval Augmented Generation

### v1

@@ Abstract
Retrieval Augmented Generation (RAG) is widely employed to ground responses to queries on domain-specific documents. But do RAG implementations leave out important information or excessively include irrelevant information? To allay these concerns, it is necessary to annotate domain-specific benchmarks to evaluate information retrieval (IR) performance, as relevance definitions vary across queries and domains. Furthermore, such benchmarks should be cost-efficiently annotated to avoid annotation selection bias. In this paper, we propose DIRAS (Domain-specific Information Retrieval Annotation with Scalability), a manual-annotation-free schema that fine-tunes open-sourced LLMs to annotate relevance labels with calibrated relevance probabilities. Extensive evaluation shows that DIRAS fine-tuned models achieve GPT-4-level performance on annotating and ranking unseen (query, document) pairs, and is helpful for real-world RAG development. [^1] We will open-source all our codes, LLM generations, and human annotations.
DIRAS: Efficient LLM-Assisted Annotation of Document Relevance
in Retrieval Augmented Generation
Jingwei Ni1,2thanks: Equal Contributions., Tobias Schimanski2footnotemark: , Meihong Lin4, Mrinmaya Sachan1, Elliott Ash1, Markus Leippold2,3 1ETH Zürich       2University of Zurich       3 Swiss Finance Institute (SFI) 4 University of Electronic Science and Technology of China {jingni, msachan, ashe}@ethz.ch, meihong_lin@uestc.edu.ch {tobias.schimanski, markus.leippold}@df.uzh.ch
@@ 1 Introduction
RAG has become a popular paradigm for NLP applications Gao et al. (2024). One core phase of RAG systems is Information Retrieval (IR), which leverages cheap retrievers to filter relevant information and thus save LLM inference costs. Given its cost-saving nature, IR might be a performance bottleneck for RAG (Chen et al., 2023b; Gao et al., 2024). Both leaving out important relevant information (low recall) as well as including excessively related but irrelevant information (low precision) may lead to severe performance down-grades (Ni et al., 2023; Cuconasu et al., 2024; Niu et al., 2024; Schimanski et al., 2024a). However, evaluation results on general-domain benchmarks (Thakur et al., 2021) may hardly indicate the IR performance on RAG systems, as the definition of relevance varies drastically across different domains and use cases (Bailey et al., 2008). See App. A for an example where the relevance judgment may vary widely with or without domain expertise. Therefore, domain-specific benchmarks need to be annotated to evaluate RAG systems in different domains.
To address this scarcity of domain-specific IR benchmarking, we propose DIRAS, a pipeline to annotate domain-specific IR data at scale (illustrated in Fig. 1). Specifically, users only need to input (1) some domain-specific queries and documents and (2) definitions for what is (ir)relevant for each query. Then, DIRAS distills relevance prediction data from GPT-4 to fine-tune open-sourced LLMs, which can then be used to annotate relevance on a large scale without extra API cost, avoiding annotation selection bias (Thakur et al., 2021) with limited expenditure.
Prior work distills open-sourced LLM-based rerankers using pairwise (Qin et al., 2024) or listwise (Sun et al., 2023b; Pradeep et al., 2023) as the ranking paradigm. However, we choose a pointwise approach (i.e., predicting a relevance score per document and then ranking according to the scores) for DIRAS, to fulfill our three Desiderata for a document relevance annotator:
D1. Efficient and Effective: When benchmarking IR of RAG systems, it is important to annotate all (query, paragraph) pairs to avoid annotation selection bias, since information thoroughness is critical for many queries (e.g., overall assessment, Ni et al., 2023). Therefore, the pointwise method is more favorable as it largely outperforms list- or pairwise method in efficiency (Sun et al., 2023a). Prior work worries that pointwise ranking is efficient but not effective due to the difficulty in calibration (Sun et al., 2023a; Qin et al., 2024). However, this method lacks empirical investigation in prior work. In our work, we observe both GPT-4 and DIRAS fine-tuned LLMs achieve very competent pointwise ranking results thanks to good calibration.
D2. Improved Leverage of the Relevance Definition: Relevance and partial relevance judgments might be subjective without domain expertise or explicit annotation guidelines (Bailey et al., 2008; Saracevic, 2008; Thomas et al., 2024; see also App. A). Therefore, DIRAS explicitly puts relevance definition into relevance prediction prompts to achieve more objective and consistent results. Compared to the list- or pairwise method, the pointwise method is provided with only one document each time. Thus, it may analyze the document along the relevance definition in better detail, especially with CoT prompting (Wei et al., 2022).
D3. Richer Predictions for RAG Requirements: Listwise or pairwise ranking algorithms only predict a relative rank for documents. With only the rank, retrieving the same number of documents (top-k) for all questions is suboptimal as it is likely that different questions have different amounts of relevant information (details in § 5). In contrast, the pointwise method predicts not only ranks but also binary relevance and calibrated relevance probabilities. Both binary labels and relevance scores allow RAG systems to retrieve the actual amount of relevant information to a question. Furthermore, the calibrated relevance probability is also helpful in automatic annotation (Ni et al., 2024), indicating which annotations are partially relevant and/or check-worthy.
To evaluate DIRAS, we conduct experiments with two datasets. First, we annotate a high-quality dataset based on ChatReport [^2] https://reports.chatclimate.ai/ (Ni et al., 2023), a real-world RAG application analyzing lengthy corporate reports. This dataset also incorporates the ideas of partial relevance and labeling uncertainty. ChatReport is representative of RAG applications that are sensitive to retrieval results, as thoroughly analyzing disclosed information is crucial for assessing what is under-addressed in the reports. Evaluation on ChatReport data shows that the pointwise ranking of DIRAS achieves very satisfactory performance, even superior to the widely-adopted listwise method (Pradeep et al., 2023) (§ 3.1). The best DIRAS fine-tuned model also achieves GPT-4-level performance in terms of relevance ranking and calibration (§ 3.2).
Second, we use a DIRAS fine-tuned model to re-annotate all (query, document) combinations (43k in total) in ClimRetrieve (Schimanski et al., 2024b), a real-world record of experts’ information seeking workflow. Experiments show that DIRAS fine-tuned models successfully understand fine-grained degree of relevance (§ 4.1) and enhance their performance through improved relevance definitions (§ 4.2). Furthermore, it mitigates IR annotation bias by identifying information ignored by experts (§ 4.3), and benchmarks IR algorithms’ target domain performance upon all 43k (query, document) pairs (§ 4.4). Finally, we propose recommendations for future RAG designs based on our takeaways (§ 5).
Collectively, our contributions include:
@@
    We propose DIRAS, a framework tuning open-sourced LLMs into efficient and effective IR annotators, taking domain expertise into account.
@@
    We annotate a high-quality IR benchmark based on ChatReport with (partial) relevance labels and uncertainty labels, based on explicit relevance definitions.
@@
    We compare DIRAS fine-tuned models with a real-world information-seeking workflow by experts, showing the model accurately understands granular relevance definitions and helps mitigate annotation selection bias.
@@
Figure 1: DIRAS pipeline. RAG-specific queries, documents as inputs; calibrated small-LLM annotators as outputs.
2 DIRAS
2.1 DIRAS Pipeline
To address the outlined desiderata (D1, D2, D3) in § 1, we design the DIRAS pipeline (illustrated in Fig. 1). This pipeline comprises the training data creation and fine-tuning of small LLMs to the calibrated annotators.
@@ 2.1 DIRAS Pipeline
Training Data Creation: We create the training data from domain-specific sources like reports. As a result, the data will be composed of a set of domain-specific (query, document) pairs. Each query comprises a question and a definition indicating what is relevant or irrelevant to the question. The relevance definition can be designed by human experts or generated by LLMs following App. B. To obtain the documents for each question/query, we employ a sampling strategy using top-k relevant documents ranked by a small dense retriever. We set a top-k threshold and sample an equal number of documents within and outside of the threshold for each question. While sampling in top-k aims at covering some relevant documents, sampling outside of top-k ensures covering the broader distribution of (query, document) pairs.
Open-Sourced LLM Fine-Tuning: Once we sampled the domain-specific (query, document) pairs, we create instruction fine-tuning (IFT) data with a SOTA generic LLM $\mathcal{M}$ and the prompt template $\mathcal{P}$ (see e.g., Fig. 2). The prompt template incorporates a (query, document) pair and an instruction to predict a binary relevance label and its confidence score. Finally, the distilled IFT data is used to fine-tune open-sourced LLMs $\mathcal{M}_{o}$ (“o” stands for “open”) to conduct binary relevance prediction with confidence.
2.2 Evaluation Metrics
Both $\mathcal{M}$’s and $\mathcal{M}_{o}$’s results should be evaluated against two types of human labels. (1) Relevance labels: whether a document is helpful for answering a query or not, disagreement between different annotators needs to be resolved to obtain the final relevance label. (2) Uncertainty labels: an annotation is uncertain, if the annotators have a strong disagreement, or the majority of them agree that the (query, document) is partially relevant. We use the following metrics for evaluation:
Binary Relevance: We compute the F1 Score of models’ binary relevance prediction using relevance labels. For RAG systems, binary relevance labels are important for deciding which documents should be passed to LLMs.
Calibration: Both $\mathcal{M}$ and $\mathcal{M}_{o}$ give confidence scores, which should calibrate the binary accuracy to indicate whether the prediction is trustworthy. We use Expected Calibration Error (ECE), Brier Score, and AUROC to measure calibration performance, following Kadavath et al. (2022) and Tian et al. (2023).
Information Retrieval: The confidence scores also give a calibrated relevance probability which can be used to rank documents for each query. To directly evaluate the ranking performance, we measure nDCG and MAP upon relevance labels.
Uncertainty: If the models understand the difficulty and uncertainty caused by partial relevance, they should have lower confidence scores on samples that humans found to be uncertain. Thus we compute average precision (AP) scores between confidence and uncertainty labels. We chose AP because it is threshold-free by summarizing F1 scores of all thresholds of relevance scores. Details of computing all metrics are in App. C.
3 Experiments on ChatReport
To the best of our knowledge, there is no existing IR dataset that (1) accompanies each query with domain-specific relevance definition; and (2) provides annotations of human uncertainty caused by partial relevance. Therefore, we annotate a dataset based on ChatReport (Ni et al., 2023) to fulfill evaluation purposes in § 2.2. ChatReport is an online RAG application for corporate climate report analyses and Q&A about the reports. [^3] See https://reports.chatclimate.ai/.
Data Sampling: We sample 31 questions and 80 reports from this application. Climate reports are sampled randomly from openly accessible user submissions [^4] See https://github.com/EdisonNi-hku/chatreport. (PDF parsing details in App. D). The questions are strategically sampled to ensure representativeness and diversity. Specifically, 11 questions are the core questions used in ChatReport, which cover essential topics of sustainability disclosure. 20 questions are selected from users’ customized questions posed to the ChatReport tool. Finally, we prompt GPT-4 to draft relevance definitions for all questions (see App. B).
Train-Test Split: We split the questions into 11 for testing and 20 for training. Similarly, we split the reports into 30 for testing and 50 for training. This ensures the evaluation on unseen questions and reports. For each question, we randomly sample 60 documents – 30 each from in the top-5 and outside the top-5 (using OpenAI text-embedding-3-small as the dense retriever). Ultimately, (query, document) pairs in training split are used to create training data with relevance label and confidence score predictions (details in § 3.1). Data points in the test split are passed to human annotation.
@@ 3 Experiments on ChatReport
Test Data Annotation: We leverage relevance definitions as the annotation guidelines. A document is relevant if and only if it fulfills the definition. The data labeling process follows two steps. First, we employ two annotators who independently annotate all test data to be either relevant, irrelevant, or partially relevant. Second, we employ a subject-matter expert in corporate climate disclosure to resolve conflicts to obtain final relevance labels. Besides relevance labels, we also obtain uncertainty labels from human annotations: Whenever there is strong disagreement (co-existence of relevance and irrelevance labels) or agreement on partial relevance (two or more annotators agree on partial relevance), the data point is labeled as uncertain. Inter-annotator agreement and other details can be found in App. E.
@@
Prompt: <question>: What is the firm’s Scope 3 emission? <question_definition>: This question is looking for information about the firm’s emission in … <paragraph>: {one text chunk from a climate report} Is <paragraph> helpful for answering <question>? Provide your best guess, and confidence score from 0 to 1. Teacher LLM $\mathcal{M}$: [Reason]: {Reason why the paragraph is (un)helpful.} [Guess]: {Yes or No.} [Confidence]: {confidence score between 0.0 and 1.0.}
@@
Figure 2: Our prompt template $\mathcal{P}$ for distilling training data from $\mathcal{M}$. It is shortened for presentation. Full $\mathcal{P}$ is in Fig. 9.
3.1 How to Distill Training Data?
@@
Point-Ask 39.27 84.07 94.41 87.57 76.33
Point-Ask-Prob-D 44.74 84.52 93.72 88.39 77.84
Point-Tok-D 28.83 86.32 93.31 80.90 72.52
Point-Ask-D 54.01 86.32 94.41 88.48 80.80
Table 1: GPT-4’s performance on ChatReport test set with different ranking methods (Point- or Listwise), with/without relevance definition (D), and calibration method (Ask or Tok). Unc. denotes the Average Precision (AP) of predicting uncertainty; Bin. denotes F1 score of binary relevance prediction; Cal. is the average of AUROC, ECE, and Brier Score; Info. is the average of nDCG and MAP; Avg. denotes the average of all metrics. Best scores of each column are bolded.
@@
To train better open-sourced LLMs $\mathcal{M}_{o}$, it is crucial to distill high-quality training data from teacher LLM $\mathcal{M}$. Specifically, we compare the following three implementation choices:
@@
Pointwise vs. Listwise: The listwise method is popular in ranking data distillation given its moderate cost and good performance (Sun et al., 2023b; Pradeep et al., 2023). However, the more efficient pointwise method is under-explored in prior work – majorly due to the concern about poor calibration (Sun et al., 2023a; Qin et al., 2024). Calibration method (Tok vs. Ask): One calibration method is to get the relevance confidence by probing the model’s generation probability of the token Yes/No when predicting a document’s relevance (Tok, Liang et al., 2023). An alternative way is directly asking LLMs to verbalize confidence score, which may work better for instruction following LLMs (Ask, Tian et al., 2023). With vs. without relevance definition: As ChatReport test data is annotated based on the relevance definition, performance should increase if the model correctly takes the in-context relevance definition into consideration.
@@
Following the takeaways of Thomas et al. (2024), we design the prompt $\mathcal{P}$ for the pointwise method with detailed task/role description, relevance definition and CoT prompting (see Fig. 2 and Fig. 9, prompt without definition in Fig. 11). We use the listwise ranking prompt from Sun et al. (2023b) and Pradeep et al. (2023) (see prompt with/without definition in Fig. 13/Fig. 12). For the pointwise method, we run one variation to test prompt sensitivity: directly asking for relevance probability instead of confidence for guess (prompt in Fig. 10). As the listwise ranking is sensitive to window/step size, we run three variations with window/step sizes of 2/1, 10/5, and 20/10. text-embedding-3-small is used for listwise methods’ initial ranking. The results are shown in Table 1. We observe that: (1) With the proper calibration method (Ask), the pointwise method outperforms the listwise method. (2) The listwise method is sensitive to window size, while the pointwise method gives more consistent performance across prompts. (3) Adding relevance definition drops the listwise performance in 2 out of 3 cases, while that improves the pointwise performance. Thus we choose pointwise to be our distillation strategy.
3.2 Open-Sourced LLM Fine-Tuning

### v2

# DIRAS: Efficient LLM Annotation of Document Relevance for Retrieval Augmented Generation

@@ Abstract
Retrieval Augmented Generation (RAG) is widely employed to ground responses to queries on domain-specific documents. But do RAG implementations leave out important information when answering queries that need an integrated analysis of information (e.g., Tell me good news in the stock market today.)? To address these concerns, RAG developers need to annotate information retrieval (IR) data for their domain of interest, which is challenging because (1) domain-specific queries usually need nuanced definitions of relevance beyond shallow semantic relevance; and (2) human or GPT-4 annotation is costly and cannot cover all (query, document) pairs (i.e., annotation selection bias), thus harming the effectiveness in evaluating IR recall. To address these challenges, we propose DIRAS (Domain-specific Information Retrieval Annotation with Scalability), a manual-annotation-free schema that fine-tunes open-sourced LLMs to consider nuanced relevance definition and annotate (partial) relevance labels with calibrated relevance scores. Extensive evaluation shows that DIRAS enables smaller (8B) LLMs to achieve GPT-4-level performance on annotating and ranking unseen (query, document) pairs, and is helpful for real-world RAG development.  [^1] All code, LLM generations, and human annotations in https://github.com/EdisonNi-hku/DIRAS.
DIRAS: Efficient LLM Annotation of Document Relevance
for Retrieval Augmented Generation
Jingwei Ni1,2thanks: Equal Contributions., Tobias Schimanski2footnotemark: , Meihong Lin4, Mrinmaya Sachan1, Elliott Ash1, Markus Leippold2,3 1ETH Zurich       2University of Zurich       3 Swiss Finance Institute (SFI) 4 University of Electronic Science and Technology of China {jingni, msachan, ashe}@ethz.ch, meihong_lin@uestc.edu.ch {tobias.schimanski, markus.leippold}@df.uzh.ch
@@ 1 Introduction
RAG has become one of the most popular paradigms for NLP applications (Gao et al., 2024). One core phase of RAG systems is Information Retrieval (IR), which leverages cheap retrievers to filter relevant information and thus save LLM inference costs. However, IR can be a performance bottleneck for RAG (Chen et al., 2023; Gao et al., 2024). Both leaving out important relevant information (low recall) as well as including excessively related but irrelevant information (low precision) may lead to severe decrease in performance (Ni et al., 2023; Cuconasu et al., 2024; Niu et al., 2024; Schimanski et al., 2024a). Furthermore, evaluation results on general-domain benchmarks (Thakur et al., 2021) may hardly indicate the IR performance on RAG systems, as the definition of relevance varies drastically across different domains and use cases (Schimanski et al., 2024b; Bailey et al., 2008). To address these concerns, Saad-Falcon et al. (2023a) propose ARES to fine-tune an in-domain LM judge to evaluate context relevance. Although showing effectiveness in evaluating RAG systems for QA datasets like HotpotQA (Yang et al., 2018), ARES does not address two significant Challenges that real-world RAGs are faced with:
Figure 1: Overview of the functionality of DIRAS taking (query, relevance definition, document) triplets as input and output a binary verdict and a well-calibrated relevance score, which is sensitive to the grey-scale of partial relevance.
C1. IR Recall: RAG systems are more generally purposed than QA models, where many user queries require an integrative information analysis. For example, Write an overview of major events in World War 2; or What is good news for the stock market today? For such integrative queries, a good IR recall is necessary for comprehensive responses. Saad-Falcon et al.’s (2023a) approach evaluates the relevance of retrieved contexts (precision), but is agnostic to other important information that the RAG retriever might leave out (recall). Besides, current RAG literature (Yan et al., 2024; Wang et al., 2024) mostly relies on QA datasets (Joshi et al., 2017; Yang et al., 2018; Dinan et al., 2019; Trivedi et al., 2022) for evaluation, where the questions are less integrative and can mostly be answered by specific facts from one or few sources. For such questions, IR precision is more important than recall. As a result, the challenges of IR recall for integrative queries are heavily under-explored.
C2. Relevance Definitions and Partial Relevance: To thoroughly gather relevant information for integrative queries, the IR model should go beyond shallow semantic relationships and consider domain-specific relevance definitions. Furthermore, domain-specific requirements and subjectivity in IR annotation create a rich gray scale of partial relevance between relevant and irrelevant (Bailey et al., 2008; Saracevic, 2008; Thomas et al., 2024; also see App. A). However, partial relevance is neglected entirely in RAG context relevance evaluation (Saad-Falcon et al., 2023a; Es et al., 2024).
As a combined solution for these challenges, we propose DIRAS, a framework for efficient and effective relevance annotation. To address C1, DIRAS distills relevance annotation ability from SOTA generic teacher LLMs to small student LLMs, which can cost-efficiently annotate broad (query, document) pairs for IR recall evaluation. For better efficiency, student LLMs conduct pointwise annotation – annotating (query, document) pairs one-by-one – which is under-explored in related work (Sun et al., 2023a; Qin et al., 2024) but achieves good performance for relevance annotation. To address C2, DIRAS student LLMs are trained to comprehend nuanced relevance definitions, thus handling queries with various requirements. The student LLMs annotate binary relevance labels with well-calibrated relevance scores. Thus, the relevance scores can be used for relevance ranking and to calibrate the annotation accuracy (Ni et al., 2024). Thereby, these continuous relevance scores also address the grayscale of partial relevance (see Fig. 1 for an illustration of DIRAS functionality).
We evaluate DIRAS in three steps. First, we annotate ChatReportRetrieve to evaluate the design decisions for making DIRAS models optimized relevance annotators (§ 3.1). The evaluation shows that the fine-tuned student ($\leq$ 8B) LLMs effectively understand nuanced relevance definitions – achieving GPT-4-level performance (§ 3.3). Second, we showcase how DIRAS assists in real-life IR annotation by re-annotating ClimRetrieve (Schimanski et al., 2024b). Results show that DIRAS student LLMs can effectively capture partial relevance, leverage improved relevance definitions, mitigate annotation selection bias, and annotate benchmarking datasets for IR algorithms (§ 4.1). Third, we re-annotate document relevance for general QA and RAG datasets, showing DIRAS’s potential in generic domains (§ 4.2). Collectively, our contributions include:
@@
    We propose DIRAS, a framework fine-tuning open-sourced LLMs into efficient and effective IR annotators, taking domain expertise into account.
@@
    We annotate ChatReportRetrieve, the first IR benchmark addressing integrative queries in RAG, and providing explicit relevance definition as annotation guidelines.
@@
    We showcase how to apply DIRAS in real-world IR annotation by re-annotating ClimRetrieve and general QA datasets.
@@
Figure 2: DIRAS pipeline. Domain-specific queries, and documents as inputs; calibrated student LLM annotators as outputs.
2 DIRAS Pipeline
The DIRAS pipeline is illustrated in Fig. 2. It comprises three steps: sampling a subset of (query, document) pairs for training data creation, obtaining relevance definitions for queries, and distilling relevance annotation ability from teacher to student LLMs.
@@ 2 DIRAS Pipeline
Sampling (Query, Document) Pairs: DIRAS takes domain-specific queries and documents as input. To sample representative (query, document) pairs as training data, it first ranks documents for each query using a small dense retriever. Then, it samples an equal number of documents within and outside of top-k (a pre-defined hyperparameter) to obtain representative documents for each query. While sampling in top-k aims at covering some relevant documents, sampling outside of top-k ensures covering the broader distribution of (query, document) pairs.
Obtaining Relevance Definitions: Each query in the sampled (query, document) pairs needs to be accompanied by an explicit definition indicating what is relevant or irrelevant to the question. The relevance definition can be generated by human experts, LLMs, or in a collaboration of both. In our experiments, we find GPT-4 generates suitable relevance definitions using the prompt in App. B.
Distilling Relevance Annotations from Teacher to Student: With the sampled (query, definition, document) triplets, DIRAS creates relevance-annotation data with a SOTA generic teacher LLM $\mathcal{M}_{t}$ and the prompt template $\mathcal{P}$ (illustrated in Fig. 3). Finally, the created data is used to fine-tune student LLMs $\mathcal{M}_{s}$, which will be used to conduct broad binary relevance annotation with confidence scores for calibration (Tian et al., 2023).
3 Optimizing Relevance Annotation
To train a highly performant student LLM $\mathcal{M}_{s}$ for relevance annotation, we proceed in three steps. First, we annotate a novel task-specific dataset for our evaluation, ChatReportRetrieve (§ 3.1). Second, we assess design choices and related work baselines to find the best-performing strategy for our relevance-annotation data creation with the teacher model $\mathcal{M}_{t}$ (§ 3.2). Third, we optimize the fine-tuning of student LLMs $\mathcal{M}_{s}$ by investigating implementation variants (§ 3.3).
3.1 ChatReportRetrieve
To evaluate the teacher LLMs’ ($\mathcal{M}_{t}$) and student LLMs’ ($\mathcal{M}_{s}$) comprehension of nuanced relevance definitions, we need to provide them with the same annotation guidelines (i.e., relevance definitions) and compare their annotation performance. To the best of our knowledge, there is no existing IR dataset that provides a nuanced relevance definition for each query. Hence, we annotate ChatReportRetrieve for our evaluation.
Data and annotation guideline preparation: ChatReportRetrieve is based on real-user integrative queries from ChatReport – a chat tool for answering climate-related questions based on corporate reports [^2] https://reports.chatclimate.ai/ (Ni et al., 2023). We sample a wide range of climate reports and representative user queries about the reports to construct ChatReportRetrieve. Then, we conduct a train-test split, making sure no test set report or query is seen in the training set. Finally, we prompt GPT-4 to draft relevance definitions for all queries. GPT-4 drafted relevance definitions show a good understanding of the climate disclosure domain, according to a domain expert’s feedback. See Appendices for details of data preprocessing (App. D) and relevance definition generation (App. B).
@@ 3.1 ChatReportRetrieve
Test Data Annotation: We leverage relevance definitions as the annotation guidelines. If and only if a document addresses the relevance definition, it is deemed as (partially) relevant. We explicitly account for partial relevance when the document addresses the periphery of the definition. The data labeling process follows two steps. First, we employ two annotators who independently annotate all test data to be either relevant, irrelevant, or partially relevant. Second, we employ a subject-matter expert in corporate climate disclosure to resolve conflicts to obtain final relevance labels. Besides relevance labels, we also obtain uncertainty labels from human annotations: Whenever there is strong disagreement (co-existence of relevance and irrelevance labels) or agreement on partial relevance (two or more annotators agree on partial relevance), the data point is labeled as uncertain. Inter-annotator agreement and other details can be found in App. E.
Evaluation Metrics: LLM predictions contain a binary relevance annotation and a confidence score. They will be evaluated against relevance or uncertainty of ChatReportRetrieve labels on four dimensions: (1) Binary Relevance: We compute the F1 Score of models’ binary relevance prediction using relevance labels. Binary relevance labels are important for deciding which documents should be passed to LLMs. (2) Calibration: Confidence scores should calibrate the binary accuracy to indicate annotation quality. We use Expected Calibration Error (ECE), Brier Score, and AUROC to measure calibration performance, following Kadavath et al. (2022) and Tian et al. (2023). (3) Information Retrieval: The confidence scores also give a calibrated relevance probability which can be used to rank documents for each query. To directly evaluate the ranking performance, we measure nDCG and MAP upon relevance labels. (4) Uncertainty: If the models understand the difficulty caused by partial relevance, they should have lower confidence scores on samples that humans found uncertain. Thus we compute average precision (AP) scores between confidence and uncertainty labels. Details of computing all metrics are in App. C.
@@
Prompt: <question>: What is the firm’s Scope 3 emission? <question_definition>: This question is looking for information about the firm’s emission in … <paragraph>: {one text chunk from a climate report} Is <paragraph> helpful for answering <question>? Provide your best guess, and confidence score from 0 to 1. Teacher LLM $\mathcal{M}_{t}$: [Reason]: {Reason why the paragraph is (un)helpful.} [Guess]: {Yes or No.} [Confidence]: {confidence score between 0.0 and 1.0.}
@@
Figure 3: Our prompt template $\mathcal{P}$ for distilling training data from $\mathcal{M}_{t}$. “[Reason]” is only used in the CoT setup. It is shortened for presentation. Full $\mathcal{P}$ is in App. Fig. 9.
3.2 Optimizing the Training Data Creation
@@
RAGAs - 68.15 - 37.13 -
ARES-0-Shot 25.48 52.63 79.35 77.67 58.79
ARES-2-Shot 17.38 3.85 63.49 44.97 32.42
ARES-4-Shot 18.16 5.13 69.51 49.68 35.62
ARES-8-Shot 16.99 28.81 66.77 48.75 40.33
ARES-16-Shot 20.65 24.20 64.76 44.85 38.61
Point-Ask 39.27 84.07 90.59 87.57 75.37
Point-Ask-Prob-D 44.74 84.52 91.31 88.39 77.24
Point-Tok-D 28.83 86.32 84.48 80.90 70.53
Point-Ask-D 54.01 86.32 91.10 88.48 80.00
Table 1: GPT-4’s performance on ChatReportRetrieve test set with different ranking methods (Point- or Listwise), RAGAs and ARES-few-shot, with/without relevance definition (D), and calibration method (Ask or Tok). Bin., Cal., Info., and Unc. stand for evaluation dimensions in § 3.1.
@@
We aim to train a highly performant student LLM $\mathcal{M}_{s}$ for relevance annotation. Thus, it is crucial to identify best-performing implementation choices that can be used to distill high-quality training data from the teacher LLM $\mathcal{M}_{t}$. Specifically, we compare the following four implementation choices:
ARES few-shot vs. relevance definitions: ARES (Saad-Falcon et al., 2023a) and DIRAS have different strategies to create target-domain training data: the former uses few-shot ICL [^3] Original ARES uses few-shot ICL to create synthetic queries instead of relevance labels, which is not applicable for ChatReportRetrieve. Thus we use their ChatGPT prompt for relevance judgement, while replacing GPT-3.5 with GPT-4. while the latter uses relevance definitions. We also include RAGAs (Es et al., 2024) relevance judgement as a baseline.
Pointwise vs. Listwise: The listwise method is popular in ranking data creation given its moderate cost and good performance (Sun et al., 2023b; Pradeep et al., 2023). However, the more efficient pointwise method is under-explored in prior work – majorly due to the concern about poor calibration (Sun et al., 2023a; Qin et al., 2024).
@@
Calibration method (Tok vs. Ask): One calibration method is to get the relevance confidence by probing the model’s generation probability of the token Yes/No when predicting a document’s relevance (Tok, Liang et al., 2023). An alternative way is directly asking LLMs to verbalize confidence score, which may work better for instruction following LLMs (Ask, Tian et al., 2023).
With vs. without relevance definition: As ChatReportRetrieve test data is annotated based on the relevance definition, performance should increase if the model correctly takes the in-context relevance definition into consideration.
@@
Following the takeaways of Thomas et al. (2024), we design the prompt $\mathcal{P}$ for the pointwise method, relevance definition and CoT prompting (see Fig. 3 and Fig. 9, prompt without definition in Fig. 11). We use the listwise ranking prompt from Sun et al. (2023b) and Pradeep et al. (2023) (see prompt with/without definition in Fig. 13/Fig. 12). For the pointwise method, we run one variation to test prompt sensitivity: directly asking for relevance probability instead of confidence for guess (prompt in Fig. 10). As the listwise ranking is sensitive to window/step size, we run three variations with window/step sizes of 2/1, 10/5, and 20/10. text-embedding-3-small is used for listwise methods’ initial ranking. ARES and RAGAs settings are from the original papers. For few-shot ICL, we keep relevant/irrelevant samples balanced to avoid bias.
Takeaways: Results in Table 1 show that: (1) Few-shot ICL fails to teach domain-specific relevance. The ICL illustrations (even balanced) seem to bias GPT-4 to underperform the zero-shot setting. (2) With the proper calibration method (Ask), the pointwise method outperforms the listwise method. (3) The listwise method is sensitive to window size, while the pointwise method gives more consistent performance across prompts. (4) Adding relevance definition drops the listwise performance in 2 out of 3 cases, while that improves the pointwise performance. Thus we choose pointwise to be our distillation strategy.
3.3 Optimizing DIRAS student LLMs